OpenStudy (anonymous):

What was the day of the week on 28th May, 2006? Methodology to do this sum?

Parth (parthkohli):

Wolfram :-P

OpenStudy (anonymous):

Mathematically how to perform it?

OpenStudy (saifoo.khan):

count them. :D

OpenStudy (saifoo.khan):

Use a calender of your mobile phone.

mathslover (mathslover):

no no there is a way lemme show u

Parth (parthkohli):

4 August ---> Saturday 3 August ---> Friday Just keep reverse counting.

Parth (parthkohli):

There's a pretty complicated way.

OpenStudy (anonymous):

Gosh! This is an aps ques!

mathslover (mathslover):

28 May, 2006 = (2005 years + Period from 1.1.2006 to 28.5.2006) Odd days in 1600 years = 0 Odd days in 400 years = 0 5 years = (4 ordinary years + 1 leap year) = (4 x 1 + 1 x 2) 6 odd days Jan. Feb. March April May (31 + 28 + 31 + 30 + 28 ) = 148 days 148 days = (21 weeks + 1 day) 1 odd day. Total number of odd days = (0 + 0 + 6 + 1) = 7 0 odd day. Given day is Sunday.

mathslover (mathslover):

well prefer this: http://www.indiabix.com/aptitude/calendar/

Parth (parthkohli):

lol ^ knew it. You must use technology for anything.

OpenStudy (anonymous):

@ParthKohli if u hav patience to read d wikipedia link. pls go ahead ;) thanku @mathslover i jus saw there n didn understand the odd days concept.

Parth (parthkohli):

I don't want to learn this method, Shrey, but you do.

mathslover (mathslover):

see we can write 2005 = 1600+400+5 right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

so first calculate odd days in 1600 that is 0 in 400 = 0 and in 5 years = 4+(1 leap year)=4+2=6

mathslover (mathslover):

now calculate odd days from 1.1.2006 to 28.05.2006 total days = 148 odd days = 1 hence total odd days :0+0+6+1=7 hence the day on 28.05.2006 will be sunday

Parth (parthkohli):

Y'know, you still have to use technology in between.

OpenStudy (anonymous):

oh alrite! But how come we know that no of odd days in 1600 and 400 is 0?? cuz..Odd days of a year X = Remainder of [No. of days in the year X / 7].

OpenStudy (anonymous):

365 mod 7 = 1

OpenStudy (anonymous):

> cal 8 2011 August 2011 Su Mo Tu We Th Fr Sa 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 > cal 8 2010 August 2010 Su Mo Tu We Th Fr Sa 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31

OpenStudy (anonymous):

Oh I see you got there already @mathslover

mathslover (mathslover):

wait @shrey

OpenStudy (anonymous):

@mathslover okies

mathslover (mathslover):

100 years = 76 ordinary years + 24 leap years = (76 x 1 + 24 x 2) odd days = 124 odd days. = (17 weeks + days) 5 odd days. Number of odd days in 100 years = 5. Number of odd days in 200 years = (5 x 2) 3 odd days. Number of odd days in 300 years = (5 x 3) 1 odd day. Number of odd days in 400 years = (5 x 4 + 1) 0 odd day.

mathslover (mathslover):

Similarly, each one of 800 years, 1200 years, 1600 years, 2000 years etc. has 0 odd days.

mathslover (mathslover):

this is the concept behind that: 21/7 = 3 ( 0 => remainder) hence 0 odd days in 400 and 1600

mathslover (mathslover):

@shrey got it?

OpenStudy (anonymous):

Methodology: And you need a reference point. You could use > cal 1 1 January 1 Su Mo Tu We Th Fr Sa 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 Not formatted well, but Jan 1, 1 was a Saturday. But you still have to worry about which years have an extra day. And you have to worry about the Julian / Gregorian calendar. You have to worry about where you are, because different countries switched in different centuries. And you have to worry about stuff like this: > cal 9 1752 September 1752 Su Mo Tu We Th Fr Sa 1 2 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30

OpenStudy (anonymous):

Methodology, advance 1 day for same date every year.

OpenStudy (anonymous):

gng thru it... tanx a ton :)

mathslover (mathslover):

ur welcome @shrey